hi my name is robert and i need help with this question thankyou, A sample of 81 was taken from a population with a standard deviation of 45 kilograms. If the sample mean was 112 kilograms, the mean of any other sample of 81 will be within what range 68% of the time?
Answer by stanbon(75887)

hi my name is robert and i need help with this question thankyou, A sample of 81 was taken from a population with a standard deviation of 45 kilograms. If the sample mean was 112 kilograms, the mean of any other sample of 81 will be within what range 68% of the time?
---
std of the sample means = 45/sqrt(81) = 5
-----
Draw a normal curve.
Shade 68% around the middle.
Find the z-scores for the limits (lower and upper) of the shaded area.
z = -1 and z = +1
-----
Find the corresponding x-bar values using x-bar = zs+u
Lower: x-bar = -1*5 + 112 = 107
Upper: x-bar = +1*5 + 112 = 117
=====================================
Cheers,
Stan H.